# Local order metrics for many-particle systems across length scales

### v1

@@ III.2.1 Poisson point process
A homogeneous Poisson point process in $\mathbb{R}^{d}$ is a many-particle system in which the position of each particle is totally independent. At unit mean number density ($\rho=1$), this process can be generated in a fixed hypercubic simulation box with volume $V$ by choosing a random number $N$ from a Poisson distribution with a mean of $\rho V=V$, and then uniformly placing $N$ points in the simulation box. Such a system with no spatial correlations has
@@ III.2.2 Equilibrium Packings
where $q=2ka$, $a_{1}=(1+2\phi)^{2}/(a-\phi)^{4}$, and $a_{2}=-(1+\phi/2)^{2}/(1-\phi)^{4}$. We substitute Eq. (26) with $\phi=0.85$ for $d=1$ and Eq. (27) $\phi=0.48$ for $d=3$ into Eq. (9) to compute their respective number variances. There is no closed-form approximation for $S(k)$ for $d=2$, so we directly sample the number variance from 1000 disk packings with $\phi=0.65$ and $N=10^{4}$ generated by a Metropolis numerical scheme [1, 2].
@@ III.2.4 Poisson cluster process
The Poisson cluster process (PCP) is a strongly clustering many-particle system with large density fluctuations on large length scales, i.e., its $S(0)$ is large, but finite, and thus is nonhyperuniform. The PCP process is generated starting from a Poisson point process with a number density $\rho_{P}$ [81]. Each of these points is treated as the center of a cluster of points where the number of points in each cluster is chosen from a Poisson distribution with mean $c$. In this work, the positions of the points around the center are chosen from an isotropic Gaussian distribution with standard deviation $r_{0}$, which can be regarded as the characteristic length scale a single cluster. This model is also known as a modified Thomas process, which is an example of a Neyman-Scott process [82, 83]. We note that, because of the strong clustering in this system, $\rho^{1/d}$ overestimates the mean nearest-neighbor distance. The pair statistics for this process in $\mathbb{R}^{d}$ are given by [82]
@@
At $k=0$, $S(0)=1+c$, so this point pattern is nonhyperuniform and super-Poissonian ($S(0)>1$) with $\sigma_{N}^{2}(R)\sim R^{d}$ scaling in the large-$R$ limit. Here, we use Eqs. (29) and (9) compute the number variance for $d=1,2$ and 3 with $r_{0}\rho^{1/d}=1$, $\rho=\rho_{p}c=1$, $\rho_{p}=0.1$, and $c=10$.
@@ III.2.5 Randomly vacated lattices
Kim and Torquato [84] showed that spatially uncorrelated point vacancies in a many-particle system has the following effect on $S(\mathbf{k})$:
@@ III.2.6 Hyposurficial point process
Hyposurficial systems are a special class of nonhyperuniform state that behave like the Poisson point process in their large-scale number variance because they lack a “surface-area” term proportional to $R^{d-1}$ in the large-$R$ expansion of the number variance, i.e., $\sigma_{N}^{2}(R)\sim AR^{d}+o(R^{d-1})$ as $R\rightarrow\infty$ where $A>0$. Hyposurficial pair statistics obey the following sum rule [9]:
@@ III.2.7 Hyperplane intersection process
The hyperplane intersection process (HIP) is a hyperfluctuating many-particle system [42], meaning its number variance scales faster than the volume of the observation window, equivalently, $\lim_{k\rightarrow 0}S(k)=\infty$ [85, 20]. Such a system is also referred to as antihyperuniform because $S(k)$ diverges to +$\infty$+ ∞ as $k\rightarrow 0$, which is diametrically opposed to hyperuniformity (i.e., $S(k)\rightarrow 0$ as $k\rightarrow 0$). This super-Poissonian and antihyperuniform many-particle system is defined as the set of intersection points of randomly and independently distributed hyperplanes [83, 86]. The pair correlation function for this process in $\mathbb{R}^{d}$ for any $d>2$ is given by [85]
@@
For any $d$, this process has $\alpha=-1$ and large-$R$ scaling $\sigma^{2}_{N}(R)\sim R^{2d-1}$, and thus, this process cannot exist for $d=1$ [43]. Here, we use Eqs. (35) and (9) to compute the number variance for $d=2$ and 3. We also note that, like PCP, HIP has a high degree of clustering and thus $\rho^{1/d}$ overestimates the mean nearest-neighbor distance.
@@ III.2.8 Hard-core antihyperuniform process
where $A$ is a positive constant and $D$ is the diameter of the hard-core interaction, both chosen such that the system is numerically realizable. Such a system has $\alpha=-1/2$, and thus, is super-Poissonian and antihyperuniform with $\sigma_{N}^{2}(R)\sim R^{d+(1/2)}$ scaling in the large-$R$ limit. This pair correlation function mimics one that corresponds to a fluid at a thermodynamic critical point, though notably not one in the Ising universality class [87]. Here, we use Eq. (37) with $A=0.1$ and $\phi=0.1$ and Eq. (9) for $d=1,2$ and 3 to compute the number variance.
@@ IV.1 Design of Number-Variance Based Order Metrics
Here, we propose the use of the scaled local number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ at sampling window radius $R$ and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$, where $R_{i,j}$ are two prescribed length scales, as order metrics for antihyperuniform, nonhyperuniform, disordered, and ordered hyperuniform systems across length scales by tracking their behavior as a function of $R$. The number variance is computed analytically where possible, and computed numerically via the direct sampling method described in Appendix A otherwise. Immediately below, we justify our choice of $v_{1}(R)$ window-volume scaling and define $\Sigma_{N}(R_{i},R_{j})$.
@@
Figure 2 shows $\sigma_{N}^{2}(R)$ for each 2D model described in Sec. III divided by their respective large-$R$ scalings. In Fig. 2(a), which depicts number variance curves for 2D hyperuniform systems, each curve rapidly plateaus indicating that they attain, on average, their large-$R$ scaling at small length scales, which was also shown by Torquato and Stillinger [9]. For standard nonhyperuniform and antihyperuniform systems, the length scale at which the large-$R$ nonhyperuniform scaling sets in can vary widely across models. In the case of disordered nearly hyperuniform point patterns, the ratio $B/A$ enables one to ascertain hyperuniform and nonhyperuniform distance-scaling regimes of the number variance $\sigma_{N}^{2}(R)$ as a function of $R$ as well as the corresponding crossover distance between the hyperuniform and nonhyperuniform regimes [46]. In Table LABEL:tab:AB, we compile values of the scaling exponent $\alpha$, volume coefficient $A$, surface area coefficient $B$, and the ratio $B/A$ for the models considered in this work. By comparing the values of $B$ in Table LABEL:tab:AB with the large-$R$ behavior in Fig. 2a and the values of $A$ in Table LABEL:tab:AB with the large-$R$ behavior in Fig. 2b, one can see that $B$ can be used to rank-order the large-$R$ degree of order/disorder in hyperuniform systems, while $A$ can likewise be used for the standard nonhyperuniform systems. Clearly, across models in a given space dimension, there is a wide range of $B/A$ values, and for a given model across space dimensions, this ratio increases. For the RVL, the defect concentration $p=A$, so when $p$ is made small $B/A$ becomes large. Thus, when $p$ is sufficiently small, RVLs can have hyperuniform $\sigma_{N}^{2}(R)$ scaling up to several orders of magnitude in $R\rho^{1/d}$, while having nonhyperuniform scaling on larger length scales. The presence of well-defined hyperuniform and nonhyperuniform variance scaling regimes makes the RVL system an excellent nontrivial example of length scale-dependent order in a system and clearly exemplifies why one must specifiy a length scale when assessing the order/disorder within a system. In particular, a global assessment of order should group the RVL with other nonhyperuniform systems, meaning one would miss, e.g., that the $\mathbb{Z}_{0.01}^{2}$ number variance curve in Fig. 5 has similar oscillations and scaling behavior to the ordered hyperuniform systems (Penrose, $\mathbb{Z}^{2},\;A_{2}$) up to approximately the $B/A$ value for $\mathbb{Z}_{0.01}^{2}$ listed in Table LABEL:tab:AB (14.35). The treatment of the number variance described above only allows one to compare large scale order within a particular asymptotic scaling class (see Sec. II.2), or, in the case of nearly nonhyperuniform systems, the transition length scale at which large-$R$ scaling sets in.
@@
We choose order metrics here based on the local number variance scaled by $v_{1}(R)$ for all systems, as opposed to scaling by their respective large-$R$ behaviors, so we can fairly compare the degree of order/disorder across length scales. For example, at small $R$, one would expect systems with hard-core interactions to have very similar short-range order. However, comparing the the small-$R$ regimes in Fig. 2a and Fig. 2b, we can see that the hard-core processes (e.g., $A_{2}$ and $\mathbb{Z}^{2}$ in Fig. 2a, RSA and equilibrium hard disks in 2b) have extremely different small-$R$ behaviors. Using $v_{1}(R)$, all curves will have $\sigma_{N}^{2}(0)/v_{1}(0)=1$, enabling us to compare the small-$R$ behavior of models across the different large-$R$ scaling classes. This scaling choice is further motivated by the success of other variance-based order metrics that use a $v_{1}(R)$ scaling [25, 26]. Moreover, with this window-volume scaling, $\sigma_{N}^{2}(R)/v_{1}(R)$ will approach +$\infty$+ ∞ as $R\rightarrow\infty$ for antihyperuniform systems, $S(0)$ (positive and bounded) as $R\rightarrow\infty$ for standard nonhyperuniform systems, and 0 as $R\rightarrow\infty$ for hyperuniform systems. Thus, with the $v_{1}(R)$ scaling, these metrics still maintain the ability to assess the degree of large-$R$ order/disorder across models.
@@
where $R_{i}<R_{j}$ are two length scales, and $\Sigma_{N}(R)$, i.e., with a single argument, implies $R_{i}=0$. In addition to reducing fluctuations, this integral metric enables us to characterize the degree of order/disorder in a point pattern over a prescribed range of length scales. Here, we focus on the single-argument function $\Sigma_{N}(R)$, which provides an integrated measure of the degree of order/disorder in a system up to a length scale $R$ and contrasts with $\sigma_{N}^{2}(R)/v_{1}(R)$, which measures the degree of order/disorder only at the length scale $R$. As above, $\Sigma_{N}(R)$ will have three large-$R$ scaling types: superlinear for antihyperuniform systems, linear for nonhyperuniform systems, and sublinear for hyperuniform systems. In Sec. V, we discuss the utility of the two-argument $\Sigma_{N}(R_{i},R_{j})$.
@@ IV.2 Application of Number Variance-Based Order Metrics for Models in $d=1,2,3$
In the following subsections, we present results for the scaled local number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ obtained analytically when possible via Eq. (9) or via direct sampling (see Sec. A) and the related integrated measure $\Sigma_{N}(R_{i},R_{j})$ for the 1D, 2D, and 3D models discussed in Sec. III. To compare different models in a particular space dimension, all distances are scaled by $\rho^{1/d}$, which is a reasonable approximation of the mean nearest-neighbor distance between particles in a system (see Sec. III for exceptions).
@@
Figure 3 shows the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ as a function of the scaled window radius $R\rho$ for each of the 1D models described in Sec. III. Changes in the $\sigma_{N}^{2}(R)/v_{1}(R)$ curves as a function of $R$ indicate it is sensitive to changes in the spatial correlations between pairs of particles the corresponding length scales. In addition, intersections between these curves show that the relative rankings of order/disorder of these systems can change across length scales. One can relate the behavior of $\sigma_{N}^{2}(R)/v_{1}(R)$ at short, intermediate, and large length scales to various characteristics of their pair statistics, for example, the greatest degree of short-range order (fastest decay of $\sigma_{N}^{2}(R)/v_{1}(R)$) is attained by systems whose particles have exclusion regions, i.e. that have $h(r)=-1$ for $0<r<r^{\prime}$, which can correspond to hard-particle systems or those with strong repulsive potentials. Slower small-$R$ decay (or growth) is associated with weaker short-range anticorrelations, i.e., $h(r)>-1$ (or short-range correlations, i.e., $h(r)>0$) in the vicinity of the origin. Oscillations at small $R$ reflect local translational order and the extent to which these oscillations extend into intermediate length scales corresponds to longer-ranged translational order. Oscillations at all $R$ indicate long range order (i.e., Bragg peaks in $S(k)$). The separation of the curves into the three scaling types reflects their respective degrees of large-$R$ order: $\sigma_{N}^{2}(R)/v_{1}(R)$ curves for antihyperuniform systems continue to increase toward +$\infty$+ ∞ at large $R$, $\sigma_{N}^{2}(R)/v_{1}(R)$ curves for nonhyperuniform systems plateau to their corresponding $S(0)$ (equivalently, $A$) values at large $R$, and $\sigma_{N}^{2}(R)/v_{1}(R)$ curves hyperuniform systems continue to decrease toward 0 at large $R$. Comparing the small- and large-$R$ behavior of, e.g., RSA and URL, one can see that the (non)hyperuniformity of a system (i.e., its long-ranged order) is not necessarily related to its degree small-$R$ order/disorder.
@@
Figure 4 shows $\Sigma_{N}(R)$ as a function of $R\rho$ for each of systems considered in Fig. 3. Here, and in subsequent sections, we consider $R\rho^{1/d}=1$ (approximately 2 mean nearest-neighbor distances) to be a “short” length scale, $R\rho^{1/d}=2.5$ (approximately 5 mean nearest-neighbor distances) to be an “intermediate” length scale, and $R\rho^{1/d}=10$ (approximately 20 mean nearest-neighbor distances) to be a “long” length scale (see vertical lines in Fig. 4). In Fig. 4, it is evident that the rank-order of the 1D point patterns is different at each of the vertical lines. As described above, the rank-order of the systems at small length scales is directly related to the small-$r$ behavior of their respective $h(r)$. A notable exception to this behavior is the HC-AHU system, whose hard-core has a radial extent significantly smaller than the other hard-core systems and is significantly more disordered by comparison. At intermediate length scales, the hard-core nonhyperuniform systems become more disordered relative to the hard-core hyperuniform systems, but are still more ordered than the non-hard-core hyperuniform systems. At large length scales, one can observe the three large-$R$ scaling classes described above: nonhyperuniform systems scale linearly, hyperuniform systems have sublinear growth and antihyperuniform systems have superlinear growth. In addition, we observe crossover between nonhyperuniform hard-core system curves and non-hard-core hyperuniform system curves, as well as between the HC-AHU curve the nonhyperuniform Poisson curve. As described in Sec. IVA, the large-$R$ ($R\rho^{1/d}=10$) rank-order here for the hyperuniform and nonhyperuniform systems match the $A$ and $B$ rank-ordering, respectively, and the different values of $B/A$ for different hard-core nonhyperuniform systems explain why these curves separate from the hard-core hyperuniform systems at different length scales (cf. Table LABEL:tab:AB).
@@
Figure 6 shows $\Sigma_{N}(R)$ as a function of $R\rho^{1/2}$ for each of systems considered in Fig. 5. As above, qualitative statements regarding the relative rankings of different types of point patterns follow from the discussion of $\Sigma_{N}(R)$ for $d=1$. The key differences in rankings from the $d=1$ results include: at small length scales, the newly introduced HIP system is the most disordered at small length scales of all systems considered. Additionally, there is a significant increase in the order of the high-$\chi$ stealthy systems due to the emergence of real-space hard-core-like character in $d\geq 2$ (cf. Figs. S5 and S16). We also find that for $d=2$ $A_{2}$ is the most ordered structure, as opposed to $\mathbb{Z}^{d}$ (cf. Fig. 4), consistent with the expected asymptotic result given their respective $B$ values. Moreover, unlike for $d=1$, we do not observe a crossover at large $R$ between the non-hard-core hyperuniform systems and hard-core nonhyperuniform systems. This difference may be a result of the small-$R$ suppression of fluctuations due to a hard core being a more dominant effect in higher space dimensions because small-$R$ behavior is weighted more strongly by $\alpha_{2}(r;R)$ than large-$R$ behavior in higher dimensions [cf. Eq. (10)]. This observation is also consistent with the increasing value of $B/A$ for nonhyperuniform systems as $d$ increases [46].

### v2

@@ III.2.1 Poisson point process
A homogeneous Poisson point process in $\mathbb{R}^{d}$ is a many-particle system in which the position of each particle is totally independent. At unit mean number density ($\rho=1$), this process can be generated in a fixed hypercubic simulation box with volume $V$ by choosing a random number $N$ from a Poisson distribution with a mean of $\rho V=V$ and then uniformly placing $N$ points in the simulation box. Such a system with no spatial correlations has
@@ III.2.2 Equilibrium Packings
where $q=2ka$, $a_{1}=(1+2\phi)^{2}/(a-\phi)^{4}$, and $a_{2}=-(1+\phi/2)^{2}/(1-\phi)^{4}$. We substitute Eq. (26) with $\phi=0.85$ for $d=1$ and Eq. (27) with $\phi=0.48$ for $d=3$ into Eq. (9) to compute their respective number variances. There is no closed-form approximation for $S(k)$ for $d=2$, so we directly sample the number variance from 1000 disk packings with $\phi=0.65$ and $N=10^{4}$ generated by a Metropolis numerical scheme [1, 2].
@@ III.2.4 Poisson cluster process
The Poisson cluster process (PCP) is a strongly clustering many-particle system with large density fluctuations on large length scales, i.e., its $S(0)$ is large, but finite, and thus is nonhyperuniform. The PCP process is generated starting from a Poisson point process with a number density $\rho_{P}$ [81]. Each of these points is treated as the center of a cluster of points where the number of points in each cluster is chosen from a Poisson distribution with mean $c$. In this work, the positions of the points around the center are chosen from an isotropic Gaussian distribution with standard deviation $r_{0}$, which can be regarded as the characteristic length scale of a single cluster. This model is also known as a modified Thomas process, which is an example of a Neyman-Scott process [82, 83]. We note that, because of the strong clustering in this system, $\rho^{1/d}$ overestimates the mean nearest-neighbor distance. The pair statistics for this process in $\mathbb{R}^{d}$ are given by [82]
@@
At $k=0$, $S(0)=1+c$, so this point pattern is nonhyperuniform and super-Poissonian [$S(0)>1$] with $\sigma_{N}^{2}(R)\sim R^{d}$ scaling in the large-$R$ limit. Here, we use Eqs. (29) and (9) to compute the number variance for $d=1,2$ and 3 with $r_{0}\rho^{1/d}=1$, $\rho=\rho_{p}c=1$, $\rho_{p}=0.1$, and $c=10$.
@@ III.2.5 Randomly vacated lattices
Kim and Torquato [84] showed that spatially uncorrelated point vacancies in a many-particle system have the following effect on $S(\mathbf{k})$:
@@ III.2.6 Hyposurficial point process
Hyposurficial systems are a special class of nonhyperuniform state that behave like the Poisson point process in their large-scale number variance because they lack a “surface-area” term proportional to $R^{d-1}$ in the large-$R$ expansion of the number variance, i.e., $\sigma_{N}^{2}(R)\sim AR^{d}+o(R^{d-1})$ as $R\rightarrow\infty$, where $A>0$. Hyposurficial pair statistics obey the following sum rule [9]:
@@ III.2.7 Hyperplane intersection process
The hyperplane intersection process (HIP) is a hyperfluctuating many-particle system [42], meaning its number variance scales faster than the volume of the observation window, equivalently, $\lim_{k\rightarrow 0}S(k)=\infty$ [85, 20]. Such a system is also referred to as antihyperuniform because $S(k)$ diverges to +$\infty$+ ∞ as $k\rightarrow 0$, which is diametrically opposed to hyperuniformity [i.e., $S(k)\rightarrow 0$ as $k\rightarrow 0$]. This super-Poissonian and antihyperuniform many-particle system is defined as the set of intersection points of randomly and independently distributed hyperplanes [83, 86]. The pair correlation function for this process in $\mathbb{R}^{d}$ for any $d>2$ is given by [85]
@@
For any $d$, this process has $\alpha=-1$ and large-$R$ scaling $\sigma^{2}_{N}(R)\sim R^{2d-1}$ and thus this process cannot exist for $d=1$ [43]. Here, we use Eqs. (35) and (9) to compute the number variance for $d=2$ and 3. We also note that, like PCP, HIP has a high degree of clustering and thus $\rho^{1/d}$ overestimates the mean nearest-neighbor distance.
@@ III.2.8 Hard-core antihyperuniform process
where $A$ is a positive constant and $D$ is the diameter of the hard-core interaction, both chosen such that the system is numerically realizable. Such a system has $\alpha=-1/2$ and thus is super-Poissonian and antihyperuniform with $\sigma_{N}^{2}(R)\sim R^{d+(1/2)}$ scaling in the large-$R$ limit. This pair correlation function mimics one that corresponds to a fluid at a thermodynamic critical point, though notably not one in the Ising universality class [87]. Here, we use Eq. (37) with $A=0.1$ and $\phi=0.1$ and Eq. (9) for $d=1,2$ and 3 to compute the number variance.
@@ IV.1 Design of Number-Variance Based Order Metrics
Here, we propose the use of the scaled local number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ at sampling window radius $R$ and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$, where $R_{i,j}$ are two prescribed length scales, as order metrics for antihyperuniform, nonhyperuniform, disordered, and ordered hyperuniform systems across length scales by tracking their behavior as a function of $R$. The number variance is computed analytically where possible, and computed numerically via the direct sampling method described in the Appendix otherwise. Immediately below, we justify our choice of $v_{1}(R)$ window-volume scaling and define $\Sigma_{N}(R_{i},R_{j})$.
@@
Figure 2 shows $\sigma_{N}^{2}(R)$ for each 2D model described in Sec. III divided by their respective large-$R$ scalings. In Fig. 2(a), which depicts number variance curves for 2D hyperuniform systems, each curve rapidly plateaus indicating that they attain, on average, their large-$R$ scaling at small length scales, which was also shown by Torquato and Stillinger [9]. For standard nonhyperuniform and antihyperuniform systems, the length scale at which the large-$R$ nonhyperuniform scaling sets in can vary widely across models. In the case of disordered nearly hyperuniform point patterns, the ratio $B/A$ enables one to ascertain hyperuniform and nonhyperuniform distance-scaling regimes of the number variance $\sigma_{N}^{2}(R)$ as a function of $R$ as well as the corresponding crossover distance between the hyperuniform and nonhyperuniform regimes [46]. In Table LABEL:tab:AB, we compile values of the scaling exponent $\alpha$, volume coefficient $A$, surface area coefficient $B$, and the ratio $B/A$ for the models considered in this work. By comparing the values of $B$ in Table LABEL:tab:AB with the large-$R$ behavior in Fig. 2 (a) and the values of $A$ in Table LABEL:tab:AB with the large-$R$ behavior in Fig. 2 (b), one can see that $B$ can be used to rank order the large-$R$ degree of order/disorder in hyperuniform systems, while $A$ can likewise be used for the standard nonhyperuniform systems. Clearly, across models in a given space dimension, there is a wide range of $B/A$ values and, for a given model across space dimensions, this ratio increases. For the RVL, the defect concentration $p=A$, so when $p$ is made small $B/A$ becomes large. Thus, when $p$ is sufficiently small, RVLs can have hyperuniform $\sigma_{N}^{2}(R)$ scaling up to several orders of magnitude in $R\rho^{1/d}$, while having nonhyperuniform scaling on larger length scales. The presence of well-defined hyperuniform and nonhyperuniform variance scaling regimes makes the RVL system an excellent nontrivial example of length scale dependent order in a system and clearly exemplifies why one must specify a length scale when assessing the order/disorder within a system. In particular, a global assessment of order should group the RVL with other nonhyperuniform systems, meaning one would miss, e.g., that the $\mathbb{Z}_{0.01}^{2}$ number variance curve in Fig. 5 has similar oscillations and scaling behavior to the ordered hyperuniform systems (Penrose, $\mathbb{Z}^{2},\;A_{2}$) up to approximately the $B/A$ value for $\mathbb{Z}_{0.01}^{2}$ listed in Table LABEL:tab:AB (14.35). The treatment of the number variance described above only allows one to compare large scale order within a particular asymptotic scaling class (see Sec. II.2) or, in the case of nearly nonhyperuniform systems, the transition length scale at which large-$R$ scaling sets in.
@@
We choose order metrics here based on the local number variance scaled by $v_{1}(R)$ for all systems, as opposed to scaling by their respective large-$R$ behaviors, so we can fairly compare the degree of order/disorder across length scales. For example, at small $R$, one would expect systems with hard-core interactions to have very similar short-range order. However, comparing the small-$R$ regimes in Fig. 2 (a) and Fig. 2 (b), we can see that the hard-core processes [e.g., $A_{2}$ and $\mathbb{Z}^{2}$ in Fig. 2 (a); RSA and equilibrium hard disks in 2 (b)] have extremely different small-$R$ behaviors. Using $v_{1}(R)$, all curves will have $\sigma_{N}^{2}(0)/v_{1}(0)=1$, enabling us to compare the small-$R$ behavior of models across the different large-$R$ scaling classes. This scaling choice is further motivated by the success of other variance-based order metrics that use a $v_{1}(R)$ scaling [25, 26]. Moreover, with this window-volume scaling, $\sigma_{N}^{2}(R)/v_{1}(R)$ will approach +$\infty$+ ∞ as $R\rightarrow\infty$ for antihyperuniform systems, $S(0)$ (positive and bounded) as $R\rightarrow\infty$ for standard nonhyperuniform systems, and 0 as $R\rightarrow\infty$ for hyperuniform systems. Thus, with the $v_{1}(R)$ scaling, these metrics still maintain the ability to assess the degree of large-$R$ order/disorder across models.
@@
where $R_{i}<R_{j}$ are two length scales and $\Sigma_{N}(R)$, i.e., with a single argument, implies $R_{i}=0$. In addition to reducing fluctuations, this integral metric enables us to characterize the degree of order/disorder in a point pattern over a prescribed range of length scales. Here, we focus on the single-argument function $\Sigma_{N}(R)$, which provides an integrated measure of the degree of order/disorder in a system up to a length scale $R$ and contrasts with $\sigma_{N}^{2}(R)/v_{1}(R)$, which measures the degree of order/disorder only at the length scale $R$. As above, $\Sigma_{N}(R)$ will have three large-$R$ scaling types: superlinear for antihyperuniform systems, linear for nonhyperuniform systems, and sublinear for hyperuniform systems. In Sec. V, we discuss the utility of the two-argument $\Sigma_{N}(R_{i},R_{j})$.
@@ IV.2 Application of Number Variance-Based Order Metrics for Models in $d=1,2,3$
In the following subsections, we present results for the scaled local number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ obtained analytically when possible via Eq. (9) or via direct sampling (see the Appendix) and the related integrated measure $\Sigma_{N}(R_{i},R_{j})$ for the 1D, 2D, and 3D models discussed in Sec. III. To compare different models in a particular space dimension, all distances are scaled by $\rho^{1/d}$, which is a reasonable approximation of the mean nearest-neighbor distance between particles in a system (see Sec. III for exceptions).
@@
Figure 3 shows the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ as a function of the scaled window radius $R\rho$ for each of the 1D models described in Sec. III. Changes in the $\sigma_{N}^{2}(R)/v_{1}(R)$ curves as a function of $R$ indicate it is sensitive to changes in the spatial correlations between pairs of particles at the corresponding length scales. In addition, intersections between these curves show that the relative rankings of order/disorder of these systems can change across length scales. One can relate the behavior of $\sigma_{N}^{2}(R)/v_{1}(R)$ at short, intermediate, and large length scales to various characteristics of their pair statistics, for example, the greatest degree of short-range order [fastest decay of $\sigma_{N}^{2}(R)/v_{1}(R)$] is attained by systems whose particles have exclusion regions, i.e. that have $h(r)=-1$ for $0<r<r^{\prime}$, which can correspond to hard-particle systems or those with strong repulsive potentials. Slower small-$R$ decay (or growth) is associated with weaker short-range anticorrelations, i.e., $h(r)>-1$ (or short-range correlations, i.e., $h(r)>0$) in the vicinity of the origin. Oscillations at small $R$ reflect local translational order and the extent to which these oscillations extend into intermediate length scales corresponds to longer-ranged translational order. Oscillations at all $R$ indicate long range order (i.e., Bragg peaks in $S(k)$). The separation of the curves into the three scaling types reflects their respective degrees of large-$R$ order: $\sigma_{N}^{2}(R)/v_{1}(R)$ curves for antihyperuniform systems continue to increase toward +$\infty$+ ∞ at large $R$, $\sigma_{N}^{2}(R)/v_{1}(R)$ curves for nonhyperuniform systems plateau to their corresponding $S(0)$ (equivalently, $A$) values at large $R$, and $\sigma_{N}^{2}(R)/v_{1}(R)$ curves hyperuniform systems continue to decrease toward 0 at large $R$. Comparing the small- and large-$R$ behavior of, e.g., RSA and URL, one can see that the (non)hyperuniformity of a system (i.e., its long-ranged order) is not necessarily related to its degree small-$R$ order/disorder.
@@
Figure 4 shows $\Sigma_{N}(R)$ as a function of $R\rho$ for each of the systems considered in Fig. 3. Here, and in subsequent sections, we consider $R\rho^{1/d}=1$ (approximately 2 mean nearest-neighbor distances) to be a “short” length scale, $R\rho^{1/d}=2.5$ (approximately 5 mean nearest-neighbor distances) to be an “intermediate” length scale, and $R\rho^{1/d}=10$ (approximately 20 mean nearest-neighbor distances) to be a “long” length scale (see vertical lines in Fig. 4). In Fig. 4, it is evident that the rank-order of the 1D point patterns is different at each of the vertical lines. As described above, the rank-order of the systems at small length scales is directly related to the small-$r$ behavior of their respective $h(r)$. A notable exception to this behavior is the HC-AHU system, whose hard-core has a radial extent significantly smaller than the other hard-core systems and is significantly more disordered by comparison. At intermediate length scales, the hard-core nonhyperuniform systems become more disordered relative to the hard-core hyperuniform systems, but are still more ordered than the non-hard-core hyperuniform systems. At large length scales, one can observe the three large-$R$ scaling classes described above: nonhyperuniform systems scale linearly, hyperuniform systems have sublinear growth, and antihyperuniform systems have superlinear growth. In addition, we observe crossover between nonhyperuniform hard-core system curves and non-hard-core hyperuniform system curves, as well as between the HC-AHU curve and the nonhyperuniform Poisson curve. As described in Sec. IV A, the large-$R$ ($R\rho^{1/d}=10$) rank order here for the hyperuniform and nonhyperuniform systems match the $A$ and $B$ rank ordering, respectively, and the different values of $B/A$ for different hard-core nonhyperuniform systems explain why these curves separate from the hard-core hyperuniform systems at different length scales (cf. Table LABEL:tab:AB).
@@
Figure 6 shows $\Sigma_{N}(R)$ as a function of $R\rho^{1/2}$ for each of systems considered in Fig. 5. As above, qualitative statements regarding the relative rankings of different types of point patterns follow from the discussion of $\Sigma_{N}(R)$ for $d=1$. The key differences in rankings from the $d=1$ results include the following: at small length scales, the newly introduced HIP system is the most disordered at small length scales of all systems considered. Additionally, there is a significant increase in the order of the high-$\chi$ stealthy systems due to the emergence of real-space hard-core-like character in $d\geq 2$ (cf. Figs. S5 and S16). We also find that for $d=2$ $A_{2}$ is the most ordered structure, as opposed to $\mathbb{Z}^{d}$ (cf. Fig. 4), consistent with the expected asymptotic result given their respective $B$ values. Moreover, unlike for $d=1$, we do not observe a crossover at large $R$ between the non-hard-core hyperuniform systems and hard-core nonhyperuniform systems. This difference may be a result of the small-$R$ suppression of fluctuations due to a hard core being a more dominant effect in higher space dimensions because small-$R$ behavior is weighted more strongly by $\alpha_{2}(r;R)$ than large-$R$ behavior in higher dimensions [cf. Eq. (10)]. This observation is also consistent with the increasing value of $B/A$ for nonhyperuniform systems as $d$ increases [46].